# A Comprehensive Review of Few-shot Action Recognition

### v1

@@ III-C Benchmarks
In this section, we briefly introduce benchmark datasets for few-shot action recognition. Statistics of the datasets and commonly used experimental settings are listed in Table I.
III-C1 UCF101
@@
The UCF101 dataset [72] is a widely recognized benchmark in the field of human action recognition. These categories span five broad types: Human-Object Interaction, Body-Motion Only, Human-Human Interaction, Playing Musical Instruments, and Sports. For few-shot action recognition, the widely used split benchmark is proposed by ARN [68], where the training, validation and test sets consist of 70, 10 and 21 classes and 9154, 1421 and 2745 videos respectively.
III-C2 HMDB51
@@
The HMDB51 dataset [73] is a prominent benchmark for human action recognition in video sequences, containing 6,766 video clips categorized into 51 distinct action classes. Each video clip in HMDB51 is collected from a diverse set of sources, including movies, public databases, and YouTube, ensuring a wide variety of scenarios, backgrounds, and camera motions. The few-shot benchmark established by ARN is structured with a division of 64 action classes and 4280 videos for training, 12 classes and 1194 videos for validation, and 24 classes and 1292 videos for testing.
III-C3 Kinetics
@@
The Kinetics dataset [8] is one of the most comprehensive and influential benchmarks for human action recognition, consisting of 306245 high-quality video clips. The dataset has been released in several versions, including Kinetics-400 [74], Kinetics-600 [75], and Kinetics-700 [76], the new version increases the number of classes and video samples to enhance the dataset’s comprehensiveness. The few-shot benchmark proposed by CMN [7] includes 100 randomly selected classes from the Kinetics dataset, with each class containing 100 examples. These 100 classes are divided into 64/12/24 non-overlapping classes with 6400/1200/2400 videos designated for the training set, validation set, and test set respectively.
III-C4 Something-Something V2
@@
The Something-Something V2 dataset [77] is a significant benchmark in the field of video understanding and human action recognition. It consists of over 220,000 video clips covering 174 action classes, making it one of the largest and most diverse datasets for action recognition. Currently, there exist two widely used few-shot benchmarks, both containing 100 action categories with 64/24/12 classes for meta-training, meta-validation, and meta-test, though differing in action categories and sample sizes. SSv2-full, proposed by OTAM [69], includes all samples from the original dataset for each class. SSv2-small, proposed by LIM [70], comprises 100 samples per class.
III-C5 EPIC-Kitchens
@@
The EPIC-Kitchens dataset [78] is a large-scale benchmark specifically designed for fine-grained action recognition in egocentric (first-person) video recordings. It comprises over 55 hours of unscripted activities captured in kitchen environments by 32 participants from diverse cultural backgrounds, resulting in nearly 40,000 action segments. These segments are annotated with detailed labels, covering 125 verb classes and 352 noun classes, which describe various cooking-related actions and objects. HyRSM [71] collected the few-shot benchmark where the training and test sets contain 60/20 action classes and 66056/10646 videos respectively.
IV Methodologies
@@
In order to mitigate the challenge of overfitting or non-convergence made by insufficient samples, early methods propose to estimate more accurate novel class distributions by generating more sample features based on a few available samples. Whereas, the majority of solutions proposed for FSAR belong to the meta-learning framework. Figure 5 offers an overview of the general framework for few-shot action recognition, involving five main stages. To begin, a meta-training/meta-test episode is created containing $N$-way $K$-shot support videos of the novel category, along with query videos. Videos are then processed to extract frame-level or video-level features. Subsequently, through the enhancement of video features, a more robust and generalized video representation is achieved. Next, the prototype representations for various categories are computed. Finally, an alignment strategy is then designed to measure the similarity between the query and prototype, leading to the classification results. In this framework, three critical issues need to be primarily considered: (1) Learning video instance representations with transferable ability to enhance the quality and discriminative power of feature embeddings, thereby capturing complex, data structures in few-shot scenarios. (2) Creating accurate action prototypes from a limited number of videos to facilitate reliable comparison and classification. (3) Optimizing distance metrics to effectively measure similarity between support and query videos in a manner that benefits few-shot learning tasks, ensuring that the learned metrics generalize well to novel classes. In the remainder of this section, we provide a review of existing approaches, categorized into four groups based on the aspects they focus on.
IV-A Generative-based methods
We first review the early methods that adopt the generative-based learning paradigm. These methods use general deep learning strategies to solve the few-shot action recognition and to mitigate the challenge of overfitting or non-convergence made by insufficient samples. A common way is to generate more sample features based on a few available samples to estimate a more accurate class distribution. However, due to the complexity of video features, generating video features faces greater challenges. While image generation has made significant progress, generating videos for training deep learning architectures remains challenging due to the high dimensionality of the output space.
@@ IV-A Generative-based methods
IV-B Video Instance Representation
Few-shot action recognition requires learning video instance representations that are spatiotemporally discriminative and can exhibit strong generalization ability to novel actions. Since few-shot learning in the image domain has made significant progress, the spatial modeling techniques employed in these methods can potentially enhance the features of video data. Nonetheless, identifying novel action concepts from complex video data presents unique challenges attributable to the inherent temporal dimension in videos. Further investigation into modeling video features with enhanced generalization capabilities is imperative. Numerous works in FSAR literature have introduced innovative approaches to extracting robust features from diverse perspectives. Previous methods mainly rely on two kinds of backbones for extracting representation embeddings from videos: 2D backbones and 3D backbones.
IV-B1 2D Backbone-based
@@ IV-B Video Instance Representation
Commonly used networks include ResNet [88], GoogLeNet [89], and ViT [90], all pre-trained on large-scale image datasets such as ImageNet [91]. For instance, AFO [92] addresses video feature extraction by pre-training GoogLeNet [89] on seen classes and projecting the output into an adaptation-oriented feature space by fusing important feature dimensions. Recently, with the widespread adoption of large-scale models, researchers [93, 94, 95] have explored the application of the visual large model [96] or multimodal large models [97] in FASR. Leveraging the rich representations and strong generalization capabilities of these large models, these methods have demonstrated significant performance improvements. In few-shot action recognition, relying solely on frame-level features extracted by 2D networks is insufficient for reliable action recognition, as it does not account for the complex temporal information in videos and the need for generalization ability. Therefore, many methods propose new techniques to learn video instance representations by enhancing frame-level features.
CNN Models. Several early works [98, 83, 99] propose to employ Convolutional Neural Networks (CNNs) to model long-term relations in videos by extracting temporal dependencies. However, these CNN-based methods face limitations such as inflexibility in modeling variable-length videos and inadequate representation of dynamic features in FSAR.
@@
GNN Models. As an effective solution for learning and inference on complex relational data, Graph Neural Network (GNN) models are designed to process data structured as graphs. Several studies [100, 101, 102, 103, 104] have adopted GNNs to address temporal modeling in few-shot action recognition. For example, Hu et al. [100] propose a dual-pooling graph neural network with hierarchical customized graph pooling layers to enhance the discriminative ability in accurately selecting representative video content. The authors construct intra-video graphs to identify the most representative frames and an inter-video graph to adaptively eliminate negative relations between videos, thereby improving overall video representations. Similarly, Feng et al. [101] utilize GNNs to analyze local-global relations in videos, generating hierarchical capsules for understanding video relationships and structural information. The approach incorporates inter-video and intra-video routing processes to filter low-discriminative capsules comprehensively, ensuring efficient representations. Additionally, Wang et al. [104] leverage GNNs to explore intra-class and inter-class relations across video pairs simultaneously, enhancing task-aware video representations. In summary, the capability of GNNs to model relational structures facilitates exploring associations between frames or samples, thereby enhancing the robustness of feature representation.
@@
Attention Strategies. The attention mechanism [105] is another effective solution for modeling temporal information, as it dynamically focuses on the most relevant parts of a sequence, thereby capturing long-range dependencies and enhancing the representation of temporal dynamics. Many FSAR methods apply attention strategies to enhance the representation of video instances. TAVs [83] adapt videos of varying lengths to preserve their temporal information and encode temporal relationships across the entire video by assigning importance scores to training parameters, facilitating a more accurate capture of temporal patterns compared to deep networks trained with limited data. As well, Wang et al. [106] introduce bidirectional attention which imitates the top-down and bottom-up attention mechanism to capture subtle action details from extracted features. The top-down attention is directly driven by few-shot action recognition tasks, capturing specific action-related features like sports equipment and critical body parts, while the bottom-up attention is supervised by predicted saliency volumes, focusing on class-agnostic saliency features like athletes and objects in motion. Inspired by the idea of neural architecture search, Cao et al. [107] propose to explore optimal architectures concerning the sequence order of temporal attention and spatial attention blocks for few-shot action recognition. The authors employ a pre-defined transformer space to incorporate both temporal and spatial attention while implementing a space-shrinking strategy along with the analysis of training losses to examine the emphasis on spatial and temporal information at different stages, aiming for superior architectures with enhanced accuracy and cost efficiency. Additionally, SloshNet [108] utilizes the attention mechanism to encode complementary global and local temporal representations. Zhang et al. [109] observe that spatial misalignment between objects is more prevalent in videos compared to temporal inconsistencies. Recognizing the significance of spatial relations, they propose an improved few-shot action recognition approach that integrates both spatial and temporal information. Specifically, they introduce a novel spatial alignment cross Transformer, which not only learns to realign spatial relations but also incorporates temporal information. These methods leverage the sequence-dependent modeling capability of the attention mechanism to focus on key spatiotemporal features in videos, thereby improving the robustness of video instance representations.
@@
In action recognition, depth, and motion information have been proven to be effective guidance signals, helping enhance contextual understanding, capturing temporal dynamics, reduces bias [110, 111]. This idea has been leveraged in few-shot action recognition as well [112, 113, 114, 115]. For example, AMeFu-Net [112] discovers the scene representation bias, which is inevitably learned by supervised video recognition methods since human actions often happen in specific scene contexts. Therefore, the authors introduce depth information as a carrier of the scene for extra visual information. Wu et al. [113] proposes to extract the task-specific motion patterns of videos from the temporal differences across consecutive frames. This approach aims to discover and enhance task-shared informative motion information, thereby facilitating improved alignment between videos within the same task. Luo et al. [114] leverage compressed domain knowledge [116] and propose a Long-Short Term Cross-Transformer for few-shot video classification where the motion vectors encode temporal cues and indicate the importance of each frame. Wang et al. [115] develop a motion-augmented long-short contrastive learning method to jointly model the global contextual information and motion dynamics. To explicitly integrate the global context into the local matching process, a long-short contrastive objective is applied to enforce frame features to predict the global context of the videos that belong to the same class. The authors take frame differences as motion information and design a motion autodecoder to explicitly extract motion features between frame representations by reconstructing pixel motions.
@@
Beyond visual information, semantic information is also commonly used to enhance video representations and improve generalization to novel actions. Several methods focus on leveraging the semantic knowledge of action concepts [117, 118]. For instance, Xia et al. [117] tackles the deficiency of semantic information by introducing external knowledge learned globally to enhance video representations learned locally within individual tasks. Specifically, the authors maintain a global meta-action bank that stores representations of atomic action snippets, which can be used to compose various complex action categories. This global meta-action bank is shared across all action types and individual tasks, enriching the video feature representations learned locally in each task through a novel single-value decomposition technique. Similarly, Shi et al. [119] leverage commonsense knowledge of actions from external resources to prompt-tune a powerful pre-trained vision-language model for few-shot classification. A large-scale corpus of language descriptions of actions are collected as text proposals, to build an action knowledge base. The collection of text proposals is done by filling in a handcraft sentence template with an external action-related corpus or by extracting action-related phrases from captions of Web instruction videos. CLIP [97] is leveraged to generate matching scores of the proposals for each frame, and the scores can be treated as action semantics with strong generalization.
@@
IV-B2 3D Backbone-based
Several methods [17, 121, 68, 83, 122] draw inspiration from widely used video classification techniques [28, 27, 29, 30, 8, 123], utilizing pre-trained 3D models as the backbone to capture short-term dependencies within videos, such as C3D [27], I3D [8], R3D [124] and VideoMAE [96]. However, relying solely on 3D models may overlook long-term temporal dependencies in videos. Moreover, the challenge persists in generalizing 3D models, which may exhibit biases towards specific patterns, to a limited number of novel class video samples.
@@
IV-B3 Discussion
The performances of different video instance representation methods are illustrated in Table II. We observe that contemporaneous 3D backbone-based methods often demonstrate superior performance compared to those based on 2D backbones. For instance, Liu et al. [125] evaluated their proposed method using both 2D and 3D backbones, observing that the performance with a 3D CNN embedding network surpassed that of a 2D CNN embedding network. This advantage arises because 3D backbones are primarily pretrained on video datasets, enabling them to effectively capture temporal dependencies between frames. However, 3D backbones entail high computational complexity, limiting their practical applicability in real-world scenarios. In contrast, 2D backbone-based methods are preferred for their flexibility and efficiency in handling spatial information, accommodating variable-length videos, and maintaining low computational complexity.
TABLE II: Results of different Video Instance Representation Learning methods.

### v2

@@ 3.3 Benchmarks
In this section, we briefly introduce benchmark datasets for few-shot action recognition. Statistics of the datasets and commonly used experimental settings are listed in Table 1.
@@
UCF101. The UCF101 dataset [125] is a widely recognized benchmark in the field of human action recognition. These categories span five broad types: Human-Object Interaction, Body-Motion Only, Human-Human Interaction, Playing Musical Instruments, and Sports. For few-shot action recognition, the widely used split benchmark is proposed by ARN [188], where the training, validation and test sets consist of 70, 10 and 21 classes and 9154, 1421 and 2745 videos respectively.
@@
HMDB51. The HMDB51 dataset [63] is a prominent benchmark for human action recognition in video sequences, containing 6,766 video clips categorized into 51 distinct action classes. Each video clip in HMDB51 is collected from a diverse set of sources, including movies, public databases, and YouTube, ensuring a wide variety of scenarios, backgrounds, and camera motions. The few-shot benchmark established by ARN is structured with a division of 64 action classes and 4280 videos for training, 12 classes and 1194 videos for validation, and 24 classes and 1292 videos for testing.
Kinetics.
@@
The Kinetics dataset [15] is one of the most comprehensive and influential benchmarks for human action recognition, consisting of 306245 high-quality video clips. The dataset has been released in several versions, including Kinetics-400 [60], Kinetics-600 [16], and Kinetics-700 [17], the new version increases the number of classes and video samples to enhance the dataset’s comprehensiveness. The few-shot benchmark proposed by CMN [196] includes 100 randomly selected classes from the Kinetics dataset, with each class containing 100 examples. These 100 classes are divided into 64/12/24 non-overlapping classes with 6400/1200/2400 videos designated for the training set, validation set, and test set respectively. Figure 4 displays the performance trends of state-of-the-art methods over the years.
@@
Something-Something V2. The Something-Something V2 dataset [38] is a significant benchmark in the field of video understanding and human action recognition. It consists of over 220,000 video clips covering 174 action classes, making it one of the largest and most diverse datasets for action recognition. Currently, there exist two widely used few-shot benchmarks, both containing 100 action categories with 64/24/12 classes for meta-training, meta-validation, and meta-test, though differing in action categories and sample sizes. SSv2-full, proposed by OTAM [13], includes all samples from the original dataset for each class. SSv2-small, proposed by LIM [193], comprises 100 samples per class.
@@
EPIC-Kitchens. The EPIC-Kitchens dataset [20] is a large-scale benchmark specifically designed for fine-grained action recognition in egocentric (first-person) video recordings. It comprises over 55 hours of unscripted activities captured in kitchen environments by 32 participants from diverse cultural backgrounds, resulting in nearly 40,000 action segments. These segments are annotated with detailed labels, covering 125 verb classes and 352 noun classes, which describe various cooking-related actions and objects. HyRSM [157] collected the few-shot benchmark where the training and test sets contain 60/20 action classes and 66056/10646 videos respectively.
4 Methodologies
@@
In order to mitigate the challenge of overfitting or non-convergence made by insufficient samples, early methods propose to estimate more accurate novel class distributions by generating more sample features based on a few available samples. Reviewing the development of few-shot action recognition, early approaches are generative-based which focus on generating additional data to mitigate sample scarcity [96, 66, 30, 174]. Whereas, the majority of solutions proposed for FSAR belong to the meta-learning framework. These methods adopt ProtoNet as their backbone framework, a metric-based meta-learning approach that simulates few-shot scenarios during training to enable rapid generalization from a few samples, rather than directly memorizing action features. In contrast to other meta-learning methods (e.g., MAML [29]), ProtoNet-based methods have lower computational costs and are easier to optimize. Furthermore, they facilitate the expansion of distance metrics to enhance spatiotemporal matching performance for FSAR.
Figure 5 offers an overview of the meta-learning framework for few-shot action recognition, involving five main stages. To begin, a meta-training/meta-test episode is created containing $N$-way $K$-shot support videos of the novel category, along with query videos. Videos are then processed to extract frame-level or video-level features. Subsequently, through the enhancement of video features, a more robust and generalized video representation is achieved. Next, the prototype representations for various categories are computed. Finally, an alignment strategy is then designed to measure the similarity between the query and prototype, leading to the classification results. In this framework, three critical issues need to be primarily considered: (1) Learning video instance representations with transferable ability to enhance the quality and discriminative power of feature embeddings, thereby capturing complex, data structures in few-shot scenarios. (2) Creating accurate action prototypes from a limited number of videos to facilitate reliable comparison and classification. (3) Optimizing distance metrics to effectively measure similarity between support and query videos in a manner that benefits few-shot learning tasks, ensuring that the learned metrics generalize well to novel classes. In the remainder of this section, we provide a review of existing approaches, categorized into four groups based on the aspects they focus on.
4.1 Generative-based methods
We first review the early methods that adopt the generative-based learning paradigm. These methods use general deep learning strategies to solve the few-shot action recognition and to mitigate the challenge of overfitting or non-convergence made by insufficient samples. Several methods [184] leverage prior knowledge of the distribution of seen classes to create pseudo samples for novel classes. Although these pseudo-samples may not fully encapsulate the class distribution, they can aid in mitigating the distribution bias caused by limited novel samples to several extent. However, due to the complexity of video features, generating video features faces greater challenges. While image generation has made significant progress, generating videos for training deep learning architectures remains challenging due to the high dimensionality of the output space.
@@ 4.1 Generative-based methods
4.2 Video Instance Representation
Few-shot action recognition requires learning video instance representations that are spatiotemporally discriminative and can exhibit strong generalization ability to novel actions. Since few-shot learning in the image domain has made significant progress, the spatial modeling techniques employed in these methods can potentially enhance the features of video data. Nonetheless, identifying novel action concepts from complex video data presents unique challenges attributable to the inherent temporal dimension in videos. Further investigation into modeling video features with enhanced generalization capabilities is imperative. Numerous works in FSAR literature have introduced innovative approaches to extract robust features from diverse perspectives. Previous methods mainly rely on two kinds of backbones for extracting representation embeddings from videos: 2D backbones and 3D backbones.
4.2.1 2D Backbone-based
@@ 4.2 Video Instance Representation
Commonly used networks include ResNet [46], GoogLeNet [128], and ViT [24], all pre-trained on large-scale image datasets such as ImageNet [21]. For instance, AFO [200] addresses video feature extraction by pre-training GoogLeNet [128] on seen classes and projecting the output into an adaptation-oriented feature space by fusing important feature dimensions. Recently, with the widespread adoption of large-scale models, researchers [159, 99, 130] have explored the application of the visual large model [135] or language-image pretrained model [108] in FASR. Leveraging the rich representations and strong generalization capabilities of these models, these methods have demonstrated significant performance improvements. In few-shot action recognition, relying solely on frame-level features extracted by 2D networks is insufficient for reliable action recognition, as it does not account for the complex temporal information in videos and the need for generalization ability. Recently, Huang et al. [52] introduce Mamba architecture into FSAR for enhancing local features from fragments of a long sub-sequence and performing implicit temporal alignment. Therefore, many methods propose new techniques to learn video instance representations by enhancing frame-level features. CNN Models. Several early works [179, 7, 147] propose to employ Convolutional Neural Networks (CNNs) to model long-term relations in videos by extracting temporal dependencies. However, these CNN-based methods face limitations such as inflexibility in modeling variable-length videos and inadequate representation of dynamic features in FSAR.
@@
GNN Models. As an effective solution for learning and inference on complex relational data, Graph Neural Network (GNN) models are designed to process data structured as graphs. Several studies [50, 27, 178, 158] have adopted GNNs to address temporal modeling in few-shot action recognition. For example, Hu et al. [50] propose a dual-pooling graph neural network with hierarchical customized graph pooling layers to enhance the discriminative ability in accurately selecting representative video content. The authors construct intra-video graphs to identify the most representative frames and an inter-video graph to adaptively eliminate negative relations between videos, thereby improving overall video representations. Similarly, Feng et al. [27] utilize GNNs to analyze local-global relations in videos, generating hierarchical capsules for understanding video relationships and structural information. The approach incorporates inter-video and intra-video routing processes to filter low-discriminative capsules comprehensively, ensuring efficient representations. Additionally, Wang et al. [158] leverage GNNs to explore intra-class and inter-class relations across video pairs simultaneously, enhancing task-aware video representations. In summary, the capability of GNNs to model relational structures facilitates exploring associations between frames or samples, thereby enhancing the robustness of feature representation.
@@
Attention Strategies. The attention mechanism [141] is another effective solution for modeling temporal information, as it dynamically focuses on the most relevant parts of a sequence, thereby capturing long-range dependencies and enhancing the representation of temporal dynamics. Many FSAR methods apply attention strategies to enhance the representation of video instances. TAVs [7] adapt videos of varying lengths to preserve their temporal information and encode temporal relationships across the entire video by assigning importance scores to training parameters, facilitating a more accurate capture of temporal patterns compared to deep networks trained with limited data. As well, Wang et al. [150] introduce bidirectional attention which imitates the top-down and bottom-up attention mechanism to capture subtle action details from extracted features. The top-down attention is directly driven by few-shot action recognition tasks, capturing specific action-related features like sports equipment and critical body parts, while the bottom-up attention is supervised by predicted saliency volumes, focusing on class-agnostic saliency features like athletes and objects in motion. Inspired by the idea of neural architecture search, Cao et al. [14] propose to explore optimal architectures concerning the sequence order of temporal attention and spatial attention blocks for few-shot action recognition. The authors employ a pre-defined transformer space to incorporate both temporal and spatial attention while implementing a space-shrinking strategy along with the analysis of training losses to examine the emphasis on spatial and temporal information at different stages, aiming for superior architectures with enhanced accuracy and cost efficiency. Additionally, SloshNet [177] utilizes the attention mechanism to encode complementary global and local temporal representations. Zhang et al. [194] observe that spatial misalignment between objects is more prevalent in videos compared to temporal inconsistencies. Recognizing the significance of spatial relations, they propose an improved few-shot action recognition approach that integrates both spatial and temporal information. Specifically, they introduce a novel spatial alignment cross Transformer, which not only learns to realign spatial relations but also incorporates temporal information. Kumar et al. [65] leverage space-time attention mechanism to assimilate the point trajectory extracted from videos, and model inter-point relationships for consolidating spatio-temporal information.
These methods leverage the sequence-dependent modeling capability of the attention mechanism to focus on key spatiotemporal features in videos, thereby improving the robustness of video instance representations.
@@
In action recognition, depth, and motion information have been proven to be effective guidance signals, helping enhance contextual understanding, capturing temporal dynamics, reduces bias [148, 26]. This idea has been leveraged in few-shot action recognition as well [31, 170, 87, 160]. For example, AMeFu-Net [31] discovers the scene representation bias, which is inevitably learned by supervised video recognition methods since human actions often happen in specific scene contexts. Therefore, the authors introduce depth information as a carrier of the scene for extra visual information. Wu et al. [170] proposes to extract the task-specific motion patterns of videos from the temporal differences across consecutive frames. This approach aims to discover and enhance task-shared informative motion information, thereby facilitating improved alignment between videos within the same task. Luo et al. [87] leverage compressed domain knowledge [10] and propose a Long-Short Term Cross-Transformer for few-shot video classification where the motion vectors encode temporal cues and indicate the importance of each frame. Wang et al. [160] develop a motion-augmented long-short contrastive learning method to jointly model the global contextual information and motion dynamics. To explicitly integrate the global context into the local matching process, a long-short contrastive objective is applied to enforce frame features to predict the global context of the videos that belong to the same class. The authors take frame differences as motion information and design a motion autoencoder to explicitly extract motion features between frame representations by reconstructing pixel motions.
@@
Beyond visual information, semantic information is also commonly used to enhance video representations and improve generalization to novel actions. Several methods focus on leveraging the semantic knowledge of action concepts [117, 172, 61]. For instance, Xia et al. [172] tackles the deficiency of semantic information by introducing external knowledge learned globally to enhance video representations learned locally within individual tasks. Specifically, the authors maintain a global meta-action bank that stores representations of atomic action snippets, which can be used to compose various complex action categories. This global meta-action bank is shared across all action types and individual tasks, enriching the video feature representations learned locally in each task through a novel single-value decomposition technique. Similarly, Shi et al. [118] leverage commonsense knowledge of actions from external resources to prompt-tune a powerful pre-trained vision-language model for few-shot classification. A large-scale corpus of language descriptions of actions are collected as text proposals, to build an action knowledge base. The collection of text proposals is done by filling in a handcraft sentence template with an external action-related corpus or by extracting action-related phrases from captions of Web instruction videos. CLIP [108] is leveraged to generate matching scores of the proposals for each frame, and the scores can be treated as action semantics with strong generalization.
@@
Finetuning Strategies. The majority of the previously mentioned methods utilize models pre-trained on image dataset as a backbone, with fine-tuning all parameters in the backbone [73]. In recent years, there have been emerging multimodal pre-trained models typified by the CLIP [108], which generally do not require extensive fine-tuning. Recently, several FSAR methods [159] adopt the CLIP model as their backbone and proposed specific fine-tuning approaches tailored to CLIP to enhance the video feature extraction capabilities for the FSAR task. For instance, several studies [159, 99, 40, 169] froze the parameters of the text encoder to preserve its generalization capabilities, while fine-tuning the parameters of the visual encoder to adapt it to the video domain. Additionally, Wang et al. [155] expand the original encoders of CLIP with light-weight extensions to transfer the pre-trained model into action recognition task while maintaining the generalization of CLIP via freezing the parameters of the image encoder. Several studies propose to adopt the adapter into the encoder in CLIP to adapt the model to the FSAR data. For example, Xing et al. [176] propose the Fine-grained Multimodal Adaptation for fine-tuning the visual encoder in CLIP. This strategy is designed to capture global motion cues enhance local temporal dynamics and highlight fine-grained semantics related to actions. Differently, Li et al. [67] propose the perceiver-based adapter to recurrently capture the sequential dynamics alongside the timeline, which could perceive the order change. Cao et al. [12] introduce the Task-Adapter into the final layers of the visual encoder while maintaining the frozen parameters of the original pre-trained model. Within each Task-Adapter, the frozen self-attention layer is repurposed to carry out task-specific self-attention across various videos within the assigned task. This process enables the capture of both unique information among classes and shared information within classes, thus facilitating task-specific adaptation and enhancing subsequent metric measurements between the query feature and support prototypes. D2ST-Adapter [101] is formulated with a dual-pathway architecture to distinctly encode spatial and temporal features. This design allows for customization with anisotropic sampling densities along spatial and temporal domains, facilitating the dedicated learning of spatial and temporal features within respective pathways.
4.2.2 3D Backbone-based
Several methods [66, 6, 188, 7, 100] draw inspiration from widely used video classification techniques [132, 56, 137, 15, 44], utilizing pre-trained 3D models as the backbone to capture short-term dependencies within videos, such as C3D [56], I3D [15], R3D [138] and VideoMAE [135]. However, relying solely on 3D models may overlook long-term temporal dependencies in videos. Moreover, the challenge persists in generalizing 3D models, which may exhibit biases towards specific patterns, to a limited number of novel class video samples.
@@
4.2.3 Discussion
The performances of different video instance representation methods are illustrated in Table 2. We observe that contemporaneous 3D backbone-based methods often demonstrate superior performance compared to those based on 2D backbones. For instance, Liu et al. [82] evaluated their proposed method using both 2D and 3D backbones, observing that the performance with a 3D CNN embedding network surpassed that of a 2D CNN embedding network. This advantage arises because 3D backbones are primarily pretrained on video datasets, enabling them to effectively capture temporal dependencies between frames. However, 3D backbones entail high computational complexity, limiting their practical applicability in real-world scenarios. In contrast, 2D backbone-based methods are preferred for their flexibility and efficiency in handling spatial information, accommodating variable-length videos, and maintaining low computational complexity.
Table 2: Results of different video instance representation learning methods.